[Brahms-l] Shift report 20040106 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Tue Jan 06 2004 - 00:03:55 EST
------------------------------------------------------------
Date: Tue Jan  6 00:03:55 2004, Shift: 16:00-24:00
Supervisor: sjs, with: 
Log book#15 pages 25 - 25

Shift Summary:
------------------------------------------------------------
Store started around 21:00. Good beam since then, 
although with the intensity falling off fairly rapidly.  
STAR called to ask our support to dump the beam around 23:45.  I agreed since the beam was falling off and we were about to change run settings in any case.  MCR subsequently called and said they would have dumped soon in any case because of the buildup of debunched beam.  
Doesn't matter: STAR owes us one...
Runs 9384 and 9385 - pedestal
Run 9389 (FFS at 12 deg, B430 setting, D3 and D4 off; MRS at
90 deg, B1050 setting) - 220K trig 6
Run 9390 (FFS @12,B306,D3 and D4 off; MRS@90, B1050) 303K trig6
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Jan 6 00:04:23 2004

This archive was generated by hypermail 2.1.8 : Tue Jan 06 2004 - 00:04:41 EST